OpenStudy (anonymous):

1. Write the slope-intercept equation of the line perpendicular to y = 5/2 x - 2, which passes through the point (0, 2). A. y = 5x + 2 B. y = -2/5x + 2 C. y = 2x + 5 D. y = 1/2x + 5 2. Write the equation y - 2 = 2(x - 3) in standard form. A. y - 2x = 4 B. 2x - y = -6 C. 2x + y = 6 D. 2x - y = 4 3 .If you are finding the line perpendicular to x + 3y = 5, what is the slope of the perpendicular line? A. -3/5 B. 1/3 C. 3 D. -3

OpenStudy (anonymous):

4. Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). A. y = 7x - 10 B. y = -1/7 x + 2 C. y = 7x - 3 D. y = -7x + 10

OpenStudy (anonymous):

Which one are you working on?

OpenStudy (anonymous):

2.

OpenStudy (turingtest):

if you're on number 2 try to get all the variable (the parts with numbers) on one side of the equals sing. keep the constant on the other.

OpenStudy (anonymous):

So would it be A ?

OpenStudy (turingtest):

no because we get y - 2 = 2(x - 3) y-2=2x-6 y-2x=-4 so which answer is this (it needs to be changed a little to get rid of the negative)

OpenStudy (anonymous):

D?

OpenStudy (turingtest):

yes, good job one problem at a time form now on though...

OpenStudy (turingtest):

from*

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

Ok, remember that if you have a line with slope = m

OpenStudy (anonymous):

A perpendicular line has slope = -1/m

OpenStudy (anonymous):

If you have a line with slope = 2, a perpendicular line has slope = -1/2

OpenStudy (anonymous):

It'll have to be either A or B. right?

OpenStudy (anonymous):

Hey, what is the 4th?

OpenStudy (anonymous):

first you need to find the slope of that line.

OpenStudy (anonymous):

oh yea we were looking at the wrong one lol it's the one below it.

OpenStudy (anonymous):

I just see 3 problems

OpenStudy (anonymous):

u see in the reply area the first one.

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

remember that two parallel lines has the same slope.

OpenStudy (anonymous):

the slope of the line y = 7x + 2 is?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

Now you can substitute the coordinates of the point (1,-3) and the slope into this equation.\[y = mx+b\]

OpenStudy (anonymous):

so you get: \[-3 = 7*1 + b\]

OpenStudy (anonymous):

solve for b and it's almost done.

OpenStudy (anonymous):

b= -10

OpenStudy (anonymous):

Good job!

OpenStudy (anonymous):

So the answer is?

OpenStudy (anonymous):

A, B, C or D?

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

You're close.

OpenStudy (anonymous):

y = mx + b

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

Right! You got it!

OpenStudy (anonymous):

Ask whatever you don't have clear yet.

OpenStudy (anonymous):

Sorry for my english, I'm not proud of it!

OpenStudy (anonymous):

No i got it. Thanksa lot for helping me. U made me get 100% on this assignment. thank you so much for ur help. I will be coming to u if i have any other question tht i need help on.

OpenStudy (anonymous):

And it's okay

OpenStudy (anonymous):

Ok, good night =)